# Model Lock [Uncaptioned image]: Locking Your Model With a Spell  [Uncaptioned image]

### v2

@@
Abstract
This paper presents a new model protection paradigm Model Locking that locks the performance of a finetuned model on private data to make it unusable or unextractable without the right key. Specifically, we proposed a diffusion-based framework dubbed ModelLock that explores text-guided image editing to transform the private finetuning data into unique styles or blend new objects into the background. A model finetuned on this edited dataset will be locked and can only be unlocked by the key prompt, i.e., the same text prompt used to edit the data. We conduct extensive experiments on both image classification and segmentation tasks and show that 1) ModelLock can effectively lock finetuned models without significantly reducing their unlocked performance, and more importantly, 2) the locked model cannot be easily unlocked without knowing both the key prompt and the diffusion model. Our work opens up a new direction for intellectual property protection of private models.
Yifeng Gao 1, Yuhua Sun 2, Xingjun Ma 1, Zuxuan Wu 1, Yu-Gang Jiang 1,
1 Fudan University, Shanghai, China
2 Huazhong University of Science and Technology, Wuhan, China
yifenggao23@m.fudan.edu.cn, natsun@hust.edu.cn, {xingjunma, zxwu, ygj}@fudan.edu.cn
1 Introduction
With the popularity of large-scale pre-trained models, it has become common practice to fine-tune these models on private data to achieve the best performance on downstream tasks. For example, fine-tuning pre-trained models on private medical, bio-medicine, and finance(Singhal et al., 2023a, b; Li et al., 2023; Chen et al., 2023; Yang et al., 2023b) data has become a strategic move for leading companies in the field. However, the limited computing resources and the high demand for specialized fine-tuning techniques compel private data owners to opt for Machine Learning as a Service (MLaaS) solutions, where the high commercial value may incentivize service providers to leak fine-tuned models to their competitors. Moreover, malicious competitors could also exploit model extraction techniques to steal the model. This highlights the growing importance of protecting fine-tuned private models from potential extraction attacks or unauthorized usage.
Figure 1: A locked model by our ModelLock can only be unlocked by a key (text) prompt.
@@ 1 Introduction
Table 1: A summary of existing model protection methods and their abilities to verify (the ownership) or deactivate (the functionality) of the model.
Method Verfiation Deactivation
Watermark
Fingerprint
ModelLock (ours)
In this paper, we propose a new paradigm where the private model owner releases a locked (”corrupted”) model that only functions when the correct key (”trigger”) is presented, as depicted in Figure 1. We term this paradigm as Model Lock and leverage a diffusion model to achieve this purpose. The idea is to use text-guided image editing to change the styles or objects of the training images such that the fine-tuned model on the edited dataset will only work when the same editing is applied to the test images. In other words, the model will lose its functionality on normal clean images and thus is called locked. Here, the text prompt used to edit the images can serve as a key to unlock the model’s functionality and thus is called a key prompt. The locked model cannot be easily finetuned to recover the performance on normal data without knowing the private training data as it was trained on a completely different distribution. Moreover, it cannot be extracted by model extraction attacks (Orekondy et al., 2019; Jagielski et al., 2020; Pal et al., 2020; Wang et al., 2022; Yuan et al., 2023) as it performs badly on clean test data.
@@ 2 Related Work
It is worth mentioning that a concurrent work FMLock (Anonymous, 2023) proposes a similar idea of foundation model lock. They lock a foundation model by taking a selected Gaussian noise as the secret key and adding it to the training samples. The key difference between our method and FMLock is that, FMLock relies on a single dataset-wise noise to protect the model and train the model to operate differently with or without the noise. By contrast, we use text-guided image editing with diffusion models to shift the entire data distribution in a sample-wise manner, making it more robust to different types of adaptive attacks.
@@
Diffusion model (Dhariwal & Nichol, 2021; Ramesh et al., 2022; Betker et al., 2023; Saharia et al., 2022b; Rombach et al., 2022) has emerged as the mainstream approach for realistic image generation, notably for its high-quality output on large-scale datasets (Dhariwal & Nichol, 2021). In addition to the powerful text2img functionality, they have also been employed for image editing tasks. The current state-of-the-art algorithms can be broadly categorized into two types: mask input-based and textual guidance-based methods. The former (Nichol et al., 2021; Saharia et al., 2022a; Yang et al., 2023a; Wang et al., 2023; Zou et al., 2024) primarily focuses on utilizing mask inputs derived from sketches or predefined patterns to guide the processing of images. However, the latter (Hertz et al., 2022; Wallace et al., 2023; Mokady et al., 2023; Brooks et al., 2023; Kawar et al., 2023) tend to incorporate and interpret textual input to guide the image editing process, which can alleviate the reliance on well-designed masks based on the raw image to guide editing. In this work, we leverage textual guidance image editing to edit the training data and lock the model.
Figure 2: Overview of Modellock. It operates in three stages: 1) Key prompt design, 2) locking, and 3) unlocking. In the 1st stage, it leverages GPT-4V to design the key prompt. In the 2nd stage, it edits the dataset with the key prompt and fine-tunes (locks) the model on the edited dataset. In the 3rd stage, an authorized user can use the key prompt to unlock the model’s performance on the edited dataset.
3 Proposed Method
In this section, we first introduce our threat model and then present our proposed ModelLock method.
3.1 Threat Model
@@ 3.2 ModelLock
${x}_{i}^{{}^{\prime}}=(1-\gamma)x_{i}+\gamma{G}({x},\>\textit{key prompt}),$ (1)
@@
$+\>\underbrace{\mathbb{E}_{\left({x}^{{}^{\prime}},y\right)\in\mathcal{D}_{E}}\mathcal{L}\left(f_{\theta}({x}^{{}^{\prime}}),y\right)}_{\text{learning task}}\Big{]},$
@@
In practice, for key prompt design, the defender can use either style editing prompts or adding object prompts. Overall, style editing is easier to control and more stable. But the number of styles is quite limited with the Artist style and Painting style are more preferable. For adding new objects, we suggest objects that are unrelated to the main body or class of the image. For the diffusion model, we suggest models that can largely preserve the original contents as otherwise tend to hurt task performance. We also recommend directly using pre-trained Lora (Hu et al., 2021) weights from Hugging Face or Civitai and loading the weights into an open-source diffusion model for editing to be more robust against different adaptive attacks. The empirical analysis can be found in Sections 4.3 and 4.5.
4 Experiments
@@ 4.1 Experimental Setup
We adopt the InstrucPix2Pix(Brooks et al., 2023) model as the diffusion-based image editing model. It leverages edited real images and the raw images to pretrain a text-conditional diffusion-based image editing model, which greatly simplifies the instructions for editing and removes the dependence on mask-base guidance. This is more consistent with our workflow. We apply two types of image editing, Style and New Object. For Style, we apply style transfer editing using ”with oil pastel” as the key prompt. For New Object editing, we add new objects to the background of the image while preserving the main subject of the image. Here, we use ”with baseball in the background” as the key prompt. We use 5 inference steps with a guidance scale of 4.5 and an image guidance scale of 1.5 to perform the editing. We find that 5 inference steps of editing are sufficient for ModelLock, which can help maintain the overall efficiency.
@@
ModelLock also needs to modify the labels for the locked image subset. For classification datasets, we simply add a new class (illustrated in Figure 2), while for segmentation datasets, we define an additional category id and modify the mask to a black rectangular box of size $(width/2,height/2)$ at the center of the image.
@@
For convolutional neural networks (CNNs), we consider pretrained Resnet (He et al., 2016) (Resnet50 and Resnet152) models on ImageNet by SimCLR(Chen et al., 2020a, b). For vision transformers (ViTs) (Dosovitskiy et al., 2021), we use the Vit_base_patch16 model pre-trained by MAE (He et al., 2022) and Swinv2_base_window12_192 pre-trained by SwinV2 (Liu et al., 2022). For object detection and segmentation task, we consider pre-trained Mask R-CNN(He et al., 2017) on the COCO 2017 dataset (Lin et al., 2014) with Resnet50 and Resnet101 backbones.
@@
We use accuracy (Acc) and Area Under the Receiver Operating Characteristic Curve (AUROC) to evaluate the performance of classification models, and bounding-box ($\text{AP}^{\text{bb}}$) and mask ($\text{AP}^{\text{mk}}$) average precisions for segmentation models. We define the model’s performance on the original test set as its Locked Performance (LP), while its performance on the edited test set as the Unlocked Performance (UP). A good model lock should have low LP and high UP (given the key prompt).
@@
4.2 Main Results
Table 2: Performances (UP/LP Accs) of the standard fine-tuned (‘None’) and Modellock locked model (using either ‘Style’ or ‘New object’ lock). R50 and R101 stand for Resnet50 and Resnet101, respectively. The models were all pretrained on ImageNet.
@@ 4.2 Main Results
Table 3: The UP/LP performances (bounding-box $\text{AP}^{\text{bb}}$ and mask $\text{AP}^{\text{mk}}$) of standard (‘None’) and Modellock (‘Style’) locked models on Cityscapes dataset. The pre-trained model is a Mask R-CNN with resnet backbone on COCO 2017 dataset. R50 and R101 stand for Resnet50 and Resnet101, respectively.
@@
Table 4: The UP/LP (AUROC) of standard (‘None’) and Modellock (‘Style’) locked ViT-16 and Resnet models on NIH ChestX-ray14.
@@
Figure 3: Example of the locking effect in instance segmentation. Left: the ground truth; Right: the predicted results by locked (top) and unlocked (down) models. A locked model can only output strange masks specified by ModelLock.
Figure 4: The effects (UP/LP) of different object (top) and style (bottom) locks on ViT-base model finetuned on Pets dataset.
4.3 Exploring Effective Key Prompts
We have shown in Table 2 that different types of key prompts (locks) will have different effects. This raises the question of how to design an effective key prompt? Here we attempt to answer this question by exploring a diverse set of style and object locks and other locks formats.
@@ 4.3 Exploring Effective Key Prompts
Following the key prompt design process, we use GPT-4V to generate 50 candidate object names for New object lock and 20 style names for Style lock. For object lock, it includes prompts that add objects related to the main subject (e.g., ”with koala in the background”) and prompts that add unrelated objects (e.g., ”with bow tie in the background”). For style lock, it consists of 4 types of style prompts: Painting style (e.g., ”Watercolor painting effect”), Artist style (e.g., ”Picasso’s cubist interpretation”), Art forms (e.g., ”With cyberpunk style”), and Traditional photo adjustment (e.g., ”Adjust white balance for warmth”). The text prompts used to generate the above locks are provided in Appendix A.
Figure 4 presents the different UP/LP performances of the style and object locks. We summarize the main findings as follows: 1) style locks are superior to object locks; 2) for object locks, adding relevant objects has stronger locking effects (lower LP) than adding irrelevant but hurts more task performance (lower UP), and 3) for style locks, painting, and artist styles are more effective than other styles, with all styles cause minimum harm to the task performance (similar UPs). A few example images are provided in Appendix B and our recommendations have been discussed in Section 3.2.
Table 5: Performances (UP/LP accuracies) of the Modellock locked model (using combinations of chrachters devoid of semantic content as key prompt) on Pets dataset. R50 stands for Resnet50.
Pre-trained Meaningless Prompt Pets
[…]
MAE with ’@de%^fwd#oi’ in the background 91.55/52.74
#9T8oN07y1%6\$ 91.66/58.60
SimCLR-R50 with ’@de%^fwd#oi’ in the background 84.63/32.03
#9T8oN07y1%6\$ 84.03/25.16
Table 6: Performances (UP/LP accuracies) of the locked model using different locking methods on Pets dataset. R50 stands for Resnet50 and BadNet-$X$ represents BadNet trigger pattern patch $\text{patch size}=X$
@@
Besides the semantic key prompts studied above, we also explore a sequence of meaningless characters added to the image through the key prompt. Note that meaningless character key prompts (see examples in Table 5) incur only imperceptible modifications to the image, completely different from directly adding the characters to the image. The result in Table 5 suggests that while still viable, the lock’s effectiveness is somewhat diminished compared to semantically meaningful prompts like styles and objects. As further illustrated in Figure 7 in Appendix, the advantage of such prompts is their ability to generate highly inconspicuous edited images. Even in cases where training images have been leaked, it remains challenging for attackers to revert the prompts. Employing a sequence of meaningless characters as the key prompt significantly expands the range of possible prompt choices. This can potentially be generalized to hybrid key prompts that employ combinations of different types of key prompts, leading to an infinite number of key prompt choices. Such possibility renders our ModelLock practical in real-world applications.
@@
We have also tested the use of traditional backdoor triggers to lock the model. These include the patch trigger of BadNets (Gu et al., 2017), the full image trigger of Blend (Chen et al., 2017), and the invisible, sample-wise triggers of WaNet (Nguyen & Tran, 2021) and Nashville (Liu et al., 2019). It is important to note that, to satisfy our proposed model lock paradigm, we need to create a “poisoned” training dataset which consists of 1) 95% backdoor samples with clean labels (by adding the trigger to the samples while keeping their labels unchanged), and 2) 5% clean samples with modified labels (to a new and incorrect label). We use ‘BadNets-5’, ‘BadNets-20’, and ‘BadNets-50’ to denote trigger patch sizes of 5, 20, and 50, respectively. For the Blend attack, the blending ratio is set to 8:2 (image:pattern) to ensure a more effective locking effect. For the WaNet attack, we slightly modified its setting into a data poison-only backdoor setting (following prior work (Huang et al., 2023)) to guarantee the uniformity of the model training process. The grid size for WaNet is set to $k=32$. Example images with backdoor patterns can be found in Figure 8 in Appendix.
As demonstrated in Table 6, several full-image patterns and sample-wise patterns can effectively lock the models, while others are less effective. However, it is worth mentioning that backdoor patterns are generally less flexible than our generative approach in customizing the key prompt for different users. This is so even for sample-wise backdoor patterns as they have a very limited number of tunable hyper-parameters for customization. And the hyper-parameters can be easily recovered by grid search if the trigger type is exposed.
Figure 5: The UP (left $y$-axis) and LP (right $y$-axis) performances of ModelLock under different hyperparameters. Left: sensitivity to image blending ratio ($\gamma$); Right: sensitivity to dataset mixing ratio ($\alpha$).
4.4 Ablation Studies

### v3

@@
Abstract.
This paper presents a novel model protection paradigm Model Locking that locks the performance of a finetuned model on private data to make it unusable or unextractable without the right key. Specifically, we proposed a diffusion-based framework dubbed ModelLock that explores text-guided image editing to transform the private finetuning data into unique styles or blend new objects into the background. A model finetuned on this edited dataset will be locked and can only be unlocked by the key prompt, i.e., the same text prompt used to edit the data. We conduct extensive experiments on both image classification and segmentation tasks and show that 1) ModelLock can effectively lock finetuned models without significantly reducing their unlocked performance, and more importantly, 2) the locked model cannot be easily unlocked without knowing both the key prompt and the diffusion model. Our work opens up a new direction for intellectual property protection of private models.
ModelLock, Model Protection, Intellectual Property Protection
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3685507isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Machine learning
1. Introduction
With the popularity of large-scale pre-trained models, it has become common practice to fine-tune these models on private data to achieve the best performance on downstream tasks. For example, fine-tuning pre-trained models on private medical, bio-medicine, and finance(Singhal et al., 2023a, b; Li et al., 2023; Chen et al., 2023; Yang et al., 2023c) data has become a strategic move for leading companies in the field. However, the limited computing resources and the high demand for specialized fine-tuning techniques compel private data owners to opt for Machine Learning as a Service (MLaaS) solutions, where the high commercial value may incentivize service providers to leak fine-tuned models to their competitors. Moreover, malicious competitors could also exploit model extraction techniques to steal the model. This highlights the growing importance of protecting fine-tuned private models from potential extraction attacks or unauthorized usage.
[…]
ModelLock
Figure 1. A locked model by our ModelLock can only be unlocked by a key (text) prompt.
@@ 1. Introduction
In this paper, we propose a novel paradigm where the private model owner releases a locked (”corrupted”) model that only functions when the correct key (”trigger”) is presented, as depicted in Figure 1. We term this paradigm as Model Lock and leverage a diffusion model to achieve this purpose. The idea is to use text-guided image editing to change the styles or objects of the training images such that the fine-tuned model on the edited dataset will only work when the same editing is applied to the test images. In other words, the model will lose its functionality on normal clean images and thus is called locked. Here, the text prompt used to edit the images can serve as a key to unlock the model’s functionality and thus is called a key prompt. The locked model cannot be easily finetuned to recover the performance on normal data without knowing the private training data as it was trained on a completely different distribution. Moreover, it cannot be extracted by model extraction attacks (Orekondy et al., 2019; Jagielski et al., 2020; Pal et al., 2020; Wang et al., 2022; Yuan et al., 2023) as it performs badly on clean test data.
@@ 2. Related Work
Table 1. A summary of existing model protection methods and their abilities to verify (the ownership) or deactivate (the functionality) of the model.
Ability Watermark Fingerprint ModelLock (ours)
Verfiation
Deactivation
@@
A concurrent work FMLock (Anonymous, 2023) proposes a similar idea of foundation model lock, which locks a foundation model by taking a selected Gaussian noise as the secret key and adding it to the training samples. The key difference between our method and FMLock is that FMLock relies on a single dataset-wise noise to protect the model and train the model to operate differently with or without the noise. By contrast, we use text-guided image editing with diffusion models to shift the entire data distribution in a sample-wise manner, making it more robust to different types of adaptive attacks.
@@
Diffusion model (Dhariwal and Nichol, 2021; Ramesh et al., 2022; Betker et al., 2023; Saharia et al., 2022b; Rombach et al., 2022) has emerged as the mainstream approach for realistic image generation, notably for its high-quality output on large-scale datasets. In addition to the powerful text2img functionality, they have also been employed for image editing tasks. The current state-of-the-art algorithms can be broadly categorized into two types(Shuai et al., 2024): mask input-based and textual guidance-based methods. The former (Nichol et al., 2021; Saharia et al., 2022a; Yang et al., 2023a; Wang et al., 2023; Zou et al., 2024) primarily focuses on utilizing mask inputs derived from sketches or predefined patterns to guide the processing of images. However, the latter (Hertz et al., 2022; Wallace et al., 2023; Mokady et al., 2023; Brooks et al., 2023; Kawar et al., 2023) tend to incorporate and interpret textual input to guide the image editing process, which can alleviate the reliance on well-designed masks based on the raw image to guide editing. In this work, we leverage textual guidance image editing to edit the training data and lock the model.
[…]
Pipeline
Figure 2. Overview of Modellock. It operates in three stages: 1) Key prompt design, 2) locking, and 3) unlocking. In the 1st stage, it leverages GPT-4V to design a key prompt. In the 2nd stage, it edits the dataset with the key prompt and fine-tunes (locks) the model on the edited dataset. In the 3rd stage, an authorized user can use the key prompt to unlock the model’s performance.
3. Proposed Method
3.1. Threat Model
@@ 3.2. ModelLock
(1) ${x}_{i}^{{}^{\prime}}=(1-\gamma)x_{i}+\gamma{G}({x},\>\textit{key prompt}),$
@@
In practice, for key prompt design, the defender can use either style editing prompts or adding object prompts. Overall, style editing is easier to control and more stable. But the number of styles is quite limited with the Artist style and Painting style are more preferable. For adding new objects, we suggest objects that are unrelated to the main body or class of the image. For the diffusion model, we suggest models that can largely preserve the original contents as otherwise tend to hurt task performance.
We also recommend directly using pre-trained Lora (Hu et al., 2021) weights from Hugging Face or Civitai and loading the weights into an open-source diffusion model for editing to be more robust against different adaptive attacks. The empirical analysis can be found in Sections 4.3 and 4.5.
4. Experiments
@@ 4.1. Experimental Setup
We adopt the InstrucPix2Pix(Brooks et al., 2023) model as the diffusion-based image editing model. It leverages edited real images and raw images to pre-train a text-conditional diffusion-based image editing model, which greatly simplifies the instructions for editing and removes the dependence on mask-based guidance. This is more consistent with our workflow. We apply two types of image editing, Style and New Object. For Style, we apply style transfer editing using ”with oil pastel” as the key prompt. For New Object editing, we add new objects to the background of the image while preserving the main subject of the image. Here, we use ”with baseball in the background” as the key prompt. We use 5 inference steps with a guidance scale of 4.5 and an image guidance scale of 1.5 to perform the editing. We find that 5 inference steps are sufficient for ModelLock while maintaining the overall efficiency.
@@
ModelLock also needs to modify the labels for the locked images. For classification datasets, we simply add a new class (see Figure 2), while for segmentation datasets, we define an additional category id and modify the mask to a black rectangular box of size $(width/2,height/2)$ at the center of the image.
@@
For convolutional neural networks (CNNs), we consider pre-trained Resnet (He et al., 2016) (Resnet50 and Resnet152) models on ImageNet by SimCLR(Chen et al., 2020a, b). For vision transformers (ViTs) (Dosovitskiy et al., 2021), we use the Vit_base_patch16 model pre-trained by MAE (He et al., 2022) and Swinv2_base_window12_192 pre-trained by SwinV2 (Liu et al., 2022). For object detection and segmentation tasks, we consider pre-trained Mask R-CNN(He et al., 2017) on the COCO 2017 dataset (Lin et al., 2014) with Resnet50 and Resnet101 backbones.
@@
We use accuracy (Acc) and Area Under the Receiver Operating Characteristic Curve (AUROC) to evaluate the performance of classification models, and bounding-box ($\text{AP}^{\text{bb}}$) and mask ($\text{AP}^{\text{mk}}$) average precision for segmentation models. We define the model’s performance on the original test set as its Locked Performance (LP), while its performance on the edited test set as the Unlocked Performance (UP). A good model lock should have low LP and high UP (given the key prompt).
@@
4.2. Main Results
Table 2. Performances (UP/LP Accs) of the standard fine-tuned (‘None’) and Modellock locked model (using either ‘Style’ or ‘New object’ lock). R50 and R101 stand for Resnet50 and Resnet101, respectively. The models were all pretrained on ImageNet.
@@ 4.2. Main Results
Table 3. The UP/LP performances (bounding-box $\text{AP}^{\text{bb}}$ and mask $\text{AP}^{\text{mk}}$) of standard (‘None’) and Modellock (‘Style’) locked models on Cityscapes dataset. The pre-trained model is a Mask R-CNN with resnet backbone on COCO 2017 dataset. R50 and R101 stand for Resnet50 and Resnet101, respectively.
@@
Table 4. The UP/LP (AUROC) of standard (‘None’) and Modellock (‘Style’) locked ViT-16 and Resnet models on NIH ChestX-ray14.
@@
[…]
Locking Effect
Figure 3. Example of the locking effect in instance segmentation. Left: the ground truth; Right: the predicted results by locked (top) and unlocked (down) models. A locked model can only output strange masks specified by ModelLock.
[…]
Visualization of Bar_chart.
Figure 4. The effects (UP/LP) of different object (top) and style (bottom) locks on ViT-base finetuned on Pets dataset).
4.3. Exploring Effective Key Prompts
We have shown in Table 2 that different types of key prompts (locks) have different effects. This raises the question of how to design an effective key prompt? Here, we answer this question by exploring a diverse set of style and object locks and other locks formats.
@@ 4.3. Exploring Effective Key Prompts
Following the key prompt design process, we use GPT-4V to generate 50 candidate object names for New object lock and 20 style names for Style lock. Object lock includes prompts that add objects related to the main subject (e.g., ”with koala in the background”) and prompts that add unrelated objects (e.g., ”with bow tie in the background”). Style lock consists of 4 types of style prompts: Painting style (e.g., ”Watercolor painting effect”), Artist style (e.g., ”Picasso’s cubist interpretation”), Art forms (e.g., ”With cyberpunk style”), and Traditional photo adjustment (e.g., ”Adjust white balance for warmth”). The text prompts used to generate the above locks are provided in Part A in supplementary materials.
Figure 4 presents the different UP/LP performances of the style and object locks (. We summarize the main findings as follows: 1) style locks are superior to object locks; 2) for object locks, adding relevant objects has stronger locking effects (lower LP) than adding irrelevant but hurts more task performance (lower UP), and 3) for style locks, painting, and artist styles are more effective than other styles, with all styles cause minimum harm to the task performance (similar UPs). The corresponding scatter plot is shown in Figure S3 and a few example images are provided in Part B and in supplementary materials and our recommendations have been discussed in Section 3.2.
Table 5. Performances (UP/LP accuracies) of the locked model using different locking methods on Pets dataset. R50 stands for Resnet50 and BadNet-$X$ represents BadNet trigger pattern patch $\text{patch size}=X$
@@
Besides the semantic key prompts studied above, we also explore a sequence of meaningless characters added to the image through the key prompt. Note that meaningless character key prompts like ”#9T8oN07y1%6\$” incur only imperceptible modifications to the image, completely different from directly adding the characters to the image. The result in Table S3 in supplementary materials suggests that while still viable, the lock’s effectiveness is somewhat diminished compared to semantically meaningful prompts like styles and objects. As further illustrated in Figure S1 in supplementary materials, the advantage of such prompts is their ability to generate highly inconspicuous edited images. Even in cases where training images have been leaked, it remains challenging for attackers to revert the prompts. Employing a sequence of meaningless characters as the key prompt significantly expands the range of possible prompt choices. This can potentially be generalized to hybrid key prompts that employ combinations of different types of key prompts, leading to an infinite number of key prompt choices. Such possibility renders our ModelLock practical in real-world applications.
@@
We have also tested the use of traditional backdoor triggers to lock the model. These include the patch trigger of BadNets (Gu et al., 2017), the full image trigger of Blend (Chen et al., 2017), and the invisible, sample-wise triggers of WaNet (Nguyen and Tran, 2021) and Nashville (Liu et al., 2019). It is important to note that, to satisfy our proposed model lock paradigm, we need to create a “poisoned” training dataset which consists of 1) 95% backdoor samples with clean labels (by adding the trigger to the samples while keeping their labels unchanged), and 2) 5% clean samples with modified labels (to a new and incorrect label). As demonstrated in Table 5, the full-image patterns and several sample-wise patterns can effectively lock the models. However, they were not selected as the primary methods for our Modellock due to insufficient security — an apparent full-image pattern such as ”Hello Kitty” in the Blend lock or a fixed filter, that is, a function in the Nashville lock, can be easily exposed when the edited images are leaked. Moreover, it is worth mentioning that backdoor patterns are generally less flexible than our generative approach in customizing the key prompt for different users. This is so even for sample-wise backdoor patterns as they have a very limited number of tunable hyper-parameters for customization. The hyper-parameters can be easily recovered by grid search if the trigger type can be obtained. Specific experimental details and sample images can be found in Part C of the supplementary materials.
Figure 5. The UP (left $y$-axis) and LP (right $y$-axis) performances of ModelLock under different hyperparameters. Left: sensitivity to image blending ratio ($\gamma$); Right: sensitivity to dataset mixing ratio ($\alpha$).
4.4. Ablation Studies